3. La Besurta

IBÓN BAIXO DE BILLAMUERTA 15 minutes

The departure point is Besurta, located at the end of the tar trail which originates at the Hospital of Benasque (Ospital de Benás).

Follow the stream that crosses the trail, as indicated in the map. Keep walking on its right side until you reach the ibón baixo (lower mountain little lake). Its altitude is 1,940 m. This walk across the fields is easy and pleasant. Appropriate for families.

IBÓN ALTO DE BILLAMUERTA 30 more minutes (but not easy)

Once you have arrived at the ibón baixo, circumvent it on its left until you get to the ravine. Keep on going uphill following the stream to get to the ibón alto (upper little lake), which is 2,060 m. high. This second part of the excursion is rather steep and not appropriate for small children. The views over the peaks del Perdiguero, la Maladeta and the Aneto are arresting.